Consider the four points ABC , , and D , on a complex plane with affixes 2 - 3i , 1/2 , 1+ 4i and 4 + 2i respectively.

a) Plot these points on complex plane

b) Calculate the affixes of vectors AB and BC

c) Determine the affix of point E such that ABCE is a parallelogram


1
Expert's answer
2021-11-08T19:43:54-0500

a)




b)


"\overrightarrow{AB}=\dfrac{1}{2}-2+(0-(-3))i"

"\overrightarrow{AB}=-\dfrac{3}{2}+3i"


"\overrightarrow{BC}=1-\dfrac{1}{2}+(4-0)i"

"\overrightarrow{BC}=\dfrac{1}{2}+4i"

c)


"\overrightarrow{AB}=-\dfrac{3}{2}+3i=\overrightarrow{EC}"

"\overrightarrow{BC}=\dfrac{1}{2}+4i=\overrightarrow{AE}"

"\overrightarrow{EC}=1-x_E+(4-y_E)i=-\dfrac{3}{2}+3i"

"x_E=\dfrac{5}{2}, y_E=1"

The affix of point E is "\dfrac{5}{2}+i."